# X-ray Emission Signatures of Neutron Star Mergers

### v1

@@ Abstract
Neutron star (NS) mergers, including both binary NS mergers and black hole - NS mergers, are multi-messenger sources detectable in both gravitational waves (GWs) and electromagnetic (EM) radiation. The expected EM emission signatures depend on the source’s progenitor, merger remnant, and observer’s line of sight (LoS). Widely discussed EM counterparts of NS mergers have been focused in the gamma-ray (in terms of short-duration gamma-ray bursts) and optical (in terms of kilonova) bands. In this paper, we show that X-ray emission carries unique post-merger signatures that are inaccessible in other EM bands and plays an important role in understanding the physics and geometry of these mergers. We consider several binary progenitor and central engine models and investigate X-ray emission signatures from the prompt phase immediately after the merger to the afterglow phase extending years later. For the prompt phase, we devise a general method for computing phenomenological X-ray light curves and spectra for structured jets viewed from any LoS, which can be applied to X-ray observations of NS mergers to constrain the geometry. The geometric constraints can in turn be used to model the afterglow and estimate a peak time and flux—to preemptively determine afterglow characteristics would be monumental for follow-up observation campaigns of future GW sources. Finally, we provide constraints on the time window for X-ray counterpart searches of NS mergers across a range of luminosity distances and detector sensitivities.
@@
1 Introduction
The first gravitational-wave (GW) detection was made in 2015 by the LIGO and Virgo collaborations (Abbott et al., 2016), validating Einstein’s theories from a century ago and ushering in the era of GW astronomy. Since then, over 200 events have been detected, with the detection rate increasing exponentially due to advancements in detector sensitivity and data analysis techniques. The first detection of a binary neutron star (BNS) merger in 2017 through both GW (Abbott et al., 2017a) and electromagnetic (EM) observations (Abbott et al., 2017b; Goldstein et al., 2017; Savchenko et al., 2017; Troja et al., 2017; Valenti et al., 2017; Covino et al., 2017; Hallinan et al., 2017; Zhang et al., 2018a) marked the dawn of observational multi-messenger astrophysics. The source is still exhibiting interesting X-ray characteristics a few years later, which lead to interesting tests to physical models (Troja et al., 2020; Hajela et al., 2022). This remains the first and only multi-messenger detection of a BNS merger.
@@ 1 Introduction
In general, a neutron star (NS) merger involves at least one NS in the merger system. These systems, including both NS-NS mergers (also called binary NS mergers or BNS mergers), and NS - black hole (BH) mergers, are ideal multi-messenger systems that generate both GW and EM signals. Leading up to a NS merger, the binary system loses orbital energy through GW radiation, causing binary separation to decrease. The final outcome and associated GW and EM signals depends on the initial parameters of the system and the NS EoS, which is poorly constrained (Lattimer, 2021; Burgio et al., 2021). For BNS mergers, the radioactive decay of r-process elements synthesized in neutron-rich material ejected from the merger can power a luminous transient known as a kilonova (Li & Paczyński, 1998; Kulkarni, 2005; Metzger et al., 2010; Kasen et al., 2013). In the case of NS-BH mergers, if the black hole is not too massive and/or is spinning rapidly, the NS, assuming a relatively stiff EoS, can be tidally disrupted outside the innermost stable circular orbit of the BH, leaving behind neutron-rich ejecta that may produce a kilonova. Otherwise, the NS will plunge directly into the BH without significant mass ejection, resulting in little or no EM counterpart. For both BNS mergers and NS-BH mergers with tidal disruption, relativistic jets can be launched along the axis normal to the merger orbital plane, which is observable as a gamma-ray burst (GRB). In the case of the BNS merger GW170817, a short GRB 170817A was detected (Goldstein et al., 2017; Savchenko et al., 2017; Zhang et al., 2018a).
@@
GRBs are the most luminous transient in the universe marking the most violent explosions that invoke catastrophic events that launch highly relativistic jets (e.g. Zhang, 2018). Phenomenologically, they are classified into two types based on duration (Kouveliotou et al., 1993): short GRBs (sGRBs) with durations less than 2 seconds and long GRBs (lGRBs) with durations greater than 2 seconds. Multi-wavelength, multi-messenger observations have shown that in general lGRBs are associated with core-collapse of massive stars, as evidenced by their association with Type Ic supernovae (Woosley & Bloom, 2006), and that sGRBs are mostly associated from BNS mergers, as evidenced by the GW170817/GRB 170817A association (Abbott et al., 2017b). However, long-duration GRBs with no associated SNe (Della Valle et al., 2006; Fynbo et al., 2006; Gehrels et al., 2006; Gal-Yam et al., 2006; Yang et al., 2015) and short-duration GRBs likely with massive star origins (Antonelli et al., 2009; Levesque et al., 2010; Xin et al., 2011; Rossi et al., 2022) led Zhang (2006) and Zhang et al. (2007) propose that physical classifications do not always align with phenomenological classifications and that one should define GRBs based on their physical origins, i.e., Type I GRBs that originate from compact-object mergers, and Type II GRBs that are associated with the core-collapse of massive stars. The need of such a classification scheme is recently strengthened with the detection of two more long-duration GRBs associated with kilonovae (Rastinejad et al., 2022; Yang et al., 2022; Levan et al., 2024; Yang et al., 2024; Sun et al., 2025) and one short-duration GRB associated with a supernova (Ahumada et al., 2021; Rossi et al., 2022). Indeed, as discussed in Zhang (2025), besides progenitor, the duration of a GRB can be also defined by central engine, emitter, and even geometry. In this paper, we adopt the more general progenitor-based classification scheme and label NS merger-associated GRBs as Type I GRBs.
@@
Traditionally, the EM counterparts of NS mergers is believed to be most efficiently searched in the gamma-ray (for short GRB emission) and optical (for kilonova emission) bands (Metzger & Berger, 2012). The broad-band afterglow emissions for both the GRB and kilonova are also expected. Zhang (2013) proposed that GW sources due to BNS mergers can be followed by bright long-duration X-ray emission if the merger product is a long-lived magnetar instead of black hole. Such an X-ray counterpart can be sGRB-less if the line of sight is off from the sGRB jet axis (see also Sun et al. (2017, 2019)). The consequence of such a massive magnetar engine would enhance the brightness of the kilonova (Yu et al., 2013) and the broadband afterglow Gao et al. (2013). Candidate sGRB-less X-ray transients that might be of a BNS merger origin have been reported (e.g. Xue et al., 2019; Bauer et al., 2017; Quirola-Vásquez et al., 2024; Ai & Zhang, 2021), so are candidate kilonovae with a putative magnetar engine (e.g. Gao et al., 2015, 2017). It is essential to further multi-messsenger observations for future BNS merger events to study their merger products and the observing geometry, and more detailed theoretical modeling is called for.
@@
In view of the advance of the X-ray capability of studying fast X-ray transients brought by the Einstein Probe (Yuan et al., 2022) and SVOM (Wei et al., 2016) missions, in this work, we focus on the X-ray band and explore how X-ray emission signatures can vary between different binary progenitors, central engine remnants, and observer viewing angles. We develop a general method to model prompt X-ray emission from various NS mergers. We present general X-ray afterglow models that can be used to guide and estimate follow-up observations based on prompt X-ray observations. The goal of this work is to provide a comprehensive framework for modeling X-ray emissions from NS mergers across observational epochs.
@@
In Section 2 we review the evolution of NS mergers starting from various progenitors, explore possible merger remnants, consider theoretical models for long-lived central engines, and examine how viewing geometry affects observational results. In Section 3 we outline observational characteristics of emission components relevant to NS mergers, including prompt, wind, and afterglow emissions. We introduce a method to model prompt X-rays and a spindown-powered X-ray wind (for the case of a long-lived magnetar central engine) for any LoS, and produce corresponding X-ray afterglow light curves. In Section 4 we present observational expectations of NS mergers from typical progenitors and their plausible central engine remnants viewed from any angle. We highlight how one can seamlessly adopt these results to observational campaigns in X-rays by providing a summarized workflow. We conclude our findings in Section 5.
2 Central Engine
@@
Observations of GRB 170817A suggest that a jet dissipation activity initiates promptly at the time of or shortly after an NS merger (Abbott et al., 2017b, c; Goldstein et al., 2017; Savchenko et al., 2017). The source of internal energy dissipation processes that power the relativistic outflow and subsequent emission is the “central engine”. Typically, either a hyper-accreting black hole or a newly formed, rapidly rotating, and highly magnetized neutron star (magnetar) is proposed as the central engine.
The BH-NS merger product is always a black hole; thus, the formation of a hyper-accreting black hole central engine depends on whether the neutron star undergoes tidal disruption before plunging into the black hole (See upper panel of Fig. 1). The BNS merger product can either be a BH or long-lived magnetar central engine, depending on the remnant mass and equation of state (See bottom panel of Fig. 1, see also Gao et al. 2016; Margalit & Metzger 2019; Ai et al. 2020). In the following sub-sections we outline several possible central engines produced by NS mergers.
2.1 Black hole
For BH-NS mergers, if the BH is not sufficiently massive to directly swallow the NS, the NS can be tidally disrupted. In this scenario, the disrupted stellar material forms an accretion disk around the BH which can power jets (Janka et al., 1999; Rezzolla et al., 2011; Liu et al., 2017; Ruiz et al., 2021; Gottlieb et al., 2023a, b; Zhu et al., 2022) through the Blandford-Znajek mechanism (Blandford & Znajek, 1977), and probably produce late-time X-ray flares (Perna et al., 2006; Proga & Zhang, 2006).
For a non-spinning BH, the approximate criterion for tidal disruption of the NS is given by:
$r_{\rm TD}\sim\bigg{(}\frac{3M_{\bullet}}{M_{*}}\bigg{)}^{1/3}R>r_{\bullet}\sim\frac{2GM_{\bullet}}{c^{2}}$ (1)
where $r_{\rm TDE}$ is the tidal disruption radius of the NS, $r_{\bullet}$ is the event horizon of the BH, $M_{\bullet}$ and $M_{*}$ are the masses of the BH and NS, respectivley, and $R$ is the radius of the NS.
For BNS mergers, immediately before merger, some material with total mass $M_{\rm ej}$ will be dynamically ejected from the system and a remnant with total mass $M_{\rm rem}$ will remain. If $M_{\rm rem}\gtrsim 1.3\,M_{\rm TOV}$ (Ai et al., 2020; Metzger et al., 2010), a BH will be promptly formed within the dynamical timescale,
@@ 2.1 Black hole
2.2 Magnetar
While many models propose a hyper-accreting black hole as the central engine powering a GRB, Duncan & Thompson (1992) and Usov (1992) were among the first to propose that a newly formed, rapidly spinning, highly magnetized neutron star—a magnetar—could power a GRB by its enormous magnetic and spin energy reservoirs. Within the context of BNS mergers, a possible massive NS engine was first proposed by Dai et al. (2006) and Gao & Fan (2006) based on the argument that the mass of the merged remnant may be smaller than the maximum NS mass give a stiff enough NS EoS.
@@ 2.2 Magnetar
Notably, some sGRBs are followed by softer extended emission (Norris & Bonnell, 2006) and a significant fraction of sGRBs exhibit “internal” X-ray plateaus—a distinct phase characterized by shallow to flat decay followed by a very steep temporal decay which cannot be interpreted within the framework of external shock models (Rowlinson et al., 2010, 2013; Lü et al., 2015). The long-duration Type I GRB 230307A displayed a distinct X-ray component that sustained when the $\gamma$-ray emission depleted, which is consistent with the spindown of a magnetar engine (Sun et al., 2025). All these features are best interpreted with a spindown magnetar, probably with implosion (collapsing to a black hole) at the end (Zhang & Mészáros, 2001; Metzger et al., 2008; Lü et al., 2015). A black hole engine (Kisaka & Ioka, 2015; Lu & Quataert, 2023; Hamidani et al., 2024) may interpret some of the less drastic plateaus but it is apparent that in order to account for the full Type I GRB data set as observed with Swift and other telescopes, a magnetar engine is demanded at least for some GRBs.
@@
If a BNS merger remnant is not too massive to promptly collapse into a BH ($M_{\rm rem}\lesssim 1.3M_{\rm TOV}$), it may evolve through short-lived phases (See bottom panel of Fig. 1).
2.2.1 Hypermassive Neutron Stars

### v2

@@ Abstract
Neutron star (NS) mergers, including both binary NS mergers and black hole-NS mergers, are multimessenger sources detectable in both gravitational wave (GW) and electromagnetic (EM) radiation. The expected EM emission signatures depend on the source’s progenitor, merger remnant, and observer’s line of sight (LoS). Widely discussed EM counterparts of NS mergers have been focused in the gamma-ray (in terms of short-duration gamma-ray bursts) and optical (in terms of kilonova) bands. In this paper, we demonstrate that X-ray emission provides a powerful and complementary probe of post-merger physics and geometry, offering diagnostic signatures across both the prompt and long-term afterglow phases. We consider several binary progenitor and central engine models and investigate X-ray emission signatures from the prompt phase immediately after the merger to the afterglow phase extending years later. For the prompt phase, we devise a general method for computing phenomenological X-ray light curves and spectra for structured jets viewed from any LoS, which can be applied to X-ray observations of NS mergers to constrain the geometry. The geometric constraints can in turn be used to model the afterglow and estimate a peak time and flux—to preemptively determine afterglow characteristics would be monumental for follow-up observation campaigns of future GW sources. Finally, we provide constraints on the time window for X-ray counterpart searches of NS mergers across a range of luminosity distances and detector sensitivities.
@@
I Introduction
The first multimessenger detection of a binary neutron star (BNS) merger in 2017 through both GW (Abbott et al., 2017a) and electromagnetic (EM) observations (Abbott et al., 2017b; Goldstein et al., 2017; Savchenko et al., 2017; Troja et al., 2017; Valenti et al., 2017; Covino et al., 2017; Hallinan et al., 2017; Zhang et al., 2018b) opened a new chapter in observational multimessenger astrophysics. The source continues to exhibit remarkable X-ray properties years after the merger, enabling stringent tests of physical models (Troja et al., 2020; Hajela et al., 2022).
@@ I Introduction
In general, a neutron star (NS) merger is defined as consisting of at least one NS in the merger system. Such systems, including both NS-NS mergers (also called binary NS mergers, or BNS mergers), and black hole (BH)-NS mergers, are ideal multimessenger systems detectable through both GW and EM signals. Leading up to an NS merger, the binary system loses orbital energy through GW radiation, causing the binary separation to decrease. The final outcome and subsequent GW and EM signals depend on the initial parameters of the system and the NS EoS, which is poorly constrained (for recent reviews, see e.g., Lattimer, 2021; Burgio et al., 2021). For BNS mergers, the radioactive decay of r-process elements synthesized in the neutron-rich material ejected during the merger can power a luminous EM transient which peaks in the optical/infrared band, termed a “kilonova” (Li and Paczyński, 1998; Kulkarni, 2005; Metzger et al., 2010; Kasen et al., 2013). In the case of BH-NS mergers, if the mass ratio between the BH and NS is not too large, and/or the BH is spinning rapidly, and/or the NS EoS is stiff, then the NS can be tidally disrupted outside the BH innermost stable circular orbit, possibly launching jets and leaving behind neutron-rich ejecta that may power a kilonova. Otherwise, the NS will plunge directly into the BH without significant mass ejection, resulting in little to no EM counterparts. For both BNS mergers and BH-NS mergers with tidal disruption, relativistic jets can be launched perpendicular to the orbital plane, which may be detected as a gamma-ray burst (GRB).
@@
GRBs are the most luminous transients in the universe and are produced during violent explosions and cataclysmic events that launch highly relativistic jets (see e.g. Zhang, 2018). Phenomenologically, they are classified into two types based on duration (Kouveliotou et al., 1993): short GRBs—with durations less than 2 seconds—and long GRBs—with durations greater than 2 seconds. Multi-wavelength and multimessenger observations have established that long-duration GRBs are typically produced during the core-collapse of massive stars, as supported by their association with Type Ic supernovae (Woosley and Bloom, 2006). In contrast, short-duration GRBs are typically associated with NS mergers, as demonstrated by the coincident detection of a GW event from a BNS merger (GW170817) and a GRB (GRB170817A) (Abbott et al., 2017b) [^1] Unlike typical short GRBs, EM observations of GRB170817A (Goldstein et al., 2017; Savchenko et al., 2017; Zhang et al., 2018b) revealed that it possessed unusually low luminosity and soft spectral characteristics. Numerical models have suggested these observations to be consistent with a narrow jet viewed off-axis (Mooley et al., 2018a; Troja et al., 2019a; Ghirlanda et al., 2019; Ryan et al., 2024), or significant cocoon emission (Gottlieb et al., 2018; Sadeh and Waxman, 2024). This work studies the former scenario and obtains consistent results.. Numerous “peculiar” GRBs—long-duration GRBs with no associated SNe (Della Valle et al., 2006; Fynbo et al., 2006; Gehrels et al., 2006; Gal-Yam et al., 2006; Yang et al., 2015) and short-duration GRBs likely with massive star origins (Antonelli et al., 2009; Levesque et al., 2010; Xin et al., 2011; Rossi et al., 2022)—led Zhang (2006) and Zhang et al. (2007) to propose that physical classifications do not always align with phenomenological ones, and that one should define GRBs based on their physical origins, i.e., Type I GRBs that originate from compact-object mergers, and Type II GRBs that are associated with the core-collapse of massive stars. The need of such a classification scheme is recently strengthened with the detection of two additional long-duration GRBs associated with kilonovae (Rastinejad et al., 2022; Yang et al., 2022; Levan et al., 2024; Yang et al., 2024; Sun et al., 2025) and one short-duration GRB associated with a supernova (Ahumada et al., 2021; Zhang et al., 2021a; Rossi et al., 2022). As noted by Zhang (2025), GRB duration is shaped not only by the progenitor system, but also by properties of the central engine, the emission site, and the system’s geometry. In this paper, we adopt the generalized classification based on progenitor systems and label GRBs associated with NS mergers “Type I GRBs”.
@@
Traditionally, the EM counterparts of NS mergers are believed to be most efficiently searched in the gamma-ray (for GRB emission) and optical (for kilonova emission) bands (Metzger and Berger, 2012). Broadband afterglow emissions for both the GRB and kilonova are also expected. Zhang (2013) proposed that GW sources due to BNS mergers can be followed by bright long-duration X-ray emission if the merger product is a long-lived magnetar instead of black hole. Such an X-ray counterpart can be detected in the absence of a short GRB if the line of sight (LoS) is far from the jet axis (see also Sun et al. (2017, 2019)). The postulated magnetar engine would enhance the brightness of the kilonova (Yu et al., 2013) and the broadband afterglow (Gao et al., 2013). Candidate GRB-less X-ray transients potentially originating from binary neutron star (BNS) mergers have been reported (e.g., Xue et al., 2019; Bauer et al., 2017; Quirola-Vásquez et al., 2024; Ai and Zhang, 2021), along with kilonova candidates possibly boosted by a magnetar engine (e.g., Gao et al., 2015, 2017). Advancing multimessenger observations of future binary neutron star merger events is vital for probing their merger remnants and observational geometries, and relies on advancements in sophisticated theoretical modeling.
@@
Motivated by recent advances in X-ray detection capabilities—especially those enabled by fast X-ray transient (FXT) searches with the Einstein Probe (Yuan et al., 2022) and SVOM (Wei et al., 2016)—in this work we focus on the X-ray band and investigate how X-ray emission signatures vary with binary progenitors, central engine remnants, and observer viewing angles. We develop a general method to model prompt X-ray emission from various NS mergers, and release the associated open-source code PromptX (Chen et al., 2025). We present general prompt X-ray light curves and spectra that can be used to guide follow-up observations. We also employ the high-performance GRB afterglow modeling toolkit VegasAfterglow (Wang et al., 2025) in generating model X-ray afterglows and investigate their consistency with prompt emission. The goal of this work is a comprehensive framework for modeling X-ray emissions from NS mergers across observational epochs.
@@
In Section II we review the evolution of NS mergers starting from various progenitors, explore possible merger remnants, consider theoretical models for long-lived central engines, and examine how viewing geometry affects observational results. In Section III we outline observational characteristics of emission signatures relevant to NS mergers, including prompt, wind, and afterglow emissions. We introduce PromptX (Chen et al., 2025), an open-source code to model prompt X-ray emission from Type I GRBs and a spindown-powered X-ray wind (for the case of a long-lived magnetar central engine). We calculate X-ray signatures for various LoS and produce corresponding X-ray afterglow light curves. In Section IV we present observational expectations of NS mergers from typical progenitors and their plausible central engine remnants viewed from any angle. We provide a summarized workflow to outline how one can seamlessly adopt these results to observational campaigns in X-rays. We conclude our findings in Section V.
II Central Engine
@@
In compact object mergers, the accompanying high-energy EM emission is typically interpreted as arising from relativistic outflows powered by a central engine, such as a hyper-accreting black hole or a newly formed, rapidly rotating, highly magnetized neutron star (a magnetar). Observations of the gamma-ray signal associated with GW170817 are consistent with this general picture. Alternative interpretations, such as emission arising from the breakout of a broad, mildly relativistic cocoon (Kasliwal et al., 2017; Gottlieb et al., 2018), remain possible.
The BH-NS merger product is always a black hole; thus, the formation of a hyper-accreting black hole central engine depends on whether the neutron star undergoes tidal disruption before plunging into the black hole (Fig. 1, upper panel). The BNS merger product can either be a BH or long-lived magnetar central engine, depending on the total remnant mass and equation of state (Fig. 1, lower panel; see also Gao et al. 2016; Margalit and Metzger 2019; Ai et al. 2020).
II.1 Black hole
For BH-NS mergers, if certain conditions are met (e.g., comparable masses, eccentricity, stiff NS EoS, etc.), the NS can be tidally disrupted. In this scenario, part of the disrupted stellar material falls back to form an accretion disk around the black hole, which can power relativistic jets (Janka et al., 1999; Rezzolla et al., 2011; Liu et al., 2017; Ruiz et al., 2021; Gottlieb et al., 2023a, b; Zhu et al., 2022) via the Blandford-Znajek mechanism (Blandford and Znajek, 1977) or neutrino-antineutrino annihilation (Zalamea and Beloborodov, 2011) and produce late-time X-ray flares (Perna et al., 2006; Proga and Zhang, 2006); the remaining material is ejected, potentially contributing to kilonova emission (Huang et al., 2018; Zhu et al., 2020; Gompertz et al., 2023; Kawaguchi et al., 2024).
For a non-spinning BH, the approximate criterion for tidal disruption of the NS is given by
$r_{\rm TD}\sim\bigg(\frac{3M_{\bullet}}{M_{*}}\bigg)^{1/3}R>r_{\rm ISCO}\sim\frac{6GM_{\bullet}}{c^{2}},$ (1)
where $r_{\rm TD}$ is the tidal disruption radius of the NS, $r_{\rm ISCO}$ is the innermost stable circular orbit of the black hole, $M_{\bullet}$ and $M_{*}$ are the masses of the BH and NS, respectively, and $R$ is the radius of the NS. For a rotating (Kerr) black hole with dimensionless spin parameter $a_{*}$, the ISCO radius depends on the spin and whether the orbit is prograde or retrograde:
$r_{\rm ISCO}(a_{*})=\frac{GM_{\bullet}}{c^{2}}\Big[3+Z_{2}\mp\sqrt{(3-Z_{1})(3+Z_{1}+2Z_{2})}\Big],$ (2)
where the upper (minus) sign corresponds to prograde orbits ($a_{*}>0$), the lower (plus) sign to retrograde orbits ($a_{*}<0$), and
$Z_{1}$ $=$ $1+(1-a_{*}^{2})^{1/3}\Big[(1+a_{*})^{1/3}+(1-a_{*})^{1/3}\Big],$ (3)
$Z_{2}$ $=$ $\sqrt{3a_{*}^{2}+Z_{1}^{2}}.$ (4)
In the maximal spin cases, $r_{\rm ISCO}$ ranges from $GM_{\bullet}/c^{2}$ for a prograde orbit ($a_{*}=+1$) to $9GM_{\bullet}/c^{2}$ for a retrograde orbit ($a_{*}=-1$). Therefore, tidal disruption of the NS is more feasible in a prograde system with a rapidly-rotating BH.
For BNS mergers, immediately before merger, some material with total mass $M_{\rm ej}$ will be dynamically ejected from the system, and a remnant with total mass $M_{\rm rem}$ will remain. If $M_{\rm rem}\gtrsim 1.3\,M_{\rm TOV}$ (Ai et al., 2020; Metzger et al., 2010), a BH will be promptly formed within the dynamical timescale,
@@ II.1 Black hole
II.2 Magnetar
While many models propose a hyper-accreting black hole as the central engine responsible for powering a GRB, Duncan and Thompson (1992) and Usov (1992) were among the first to propose that a newly formed, rapidly spinning, highly magnetized neutron star—a magnetar—could power a GRB by its enormous magnetic and spin energy reservoirs. In the context of BNS mergers, Dai et al. (2006) and Gao and Fan (2006) proposed that a supramassive, differentially rotating millisecond magnetar remnant could power extended X-ray activity in some Type I GRBs, contingent on a sufficiently stiff NS EoS. In regards to the GRB prompt emission itself, however, early theoretical and numerical studies have demonstrated that baryon loading in the polar regions severely limits the formation of clean, ultra-relativistic jets (Murguia-Berthier et al., 2014; Ciolfi et al., 2017). However, more recent GRMHD simulations indicate that hypermassive neutron stars (discussed further in Section II.2.1) can drive mildly relativistic, magnetically dominated polar outflows, which may be further accelerated to larger Lorentz factors compatible with GRBs, suggesting a potential pathway to overcome baryon-loading limitations (Mösta et al., 2020; Most and Quataert, 2023; Musolino et al., 2025).
@@ II.2 Magnetar
Notably, some Type I GRBs are followed by softer extended emission (Norris and Bonnell, 2006), and a significant fraction of Type I GRBs exhibit “internal” X-ray plateaus—a distinct phase characterized by shallow to flat decay followed by a very steep temporal decay which cannot be interpreted within the framework of external shock models (Rowlinson et al., 2010, 2013). These features can instead be naturally explained by continued energy injection from the spindown of a magnetar engine, with the sudden decay signaling its collapse to a BH (Troja et al., 2007; Metzger et al., 2008; Zhang, 2014; Lü et al., 2015; Siegel and Ciolfi, 2016a, b; Sun et al., 2025). Interestingly, kilonova-associated GRBs appear to populate a systematically dimmer region of the X-ray plateau luminosity–duration plane (Dainotti et al., 2020). Some theoretical studies have explored possible explanations for X-ray plateaus in short GRBs that do not involve a magnetar engine (Kisaka and Ioka, 2015; Lu and Quataert, 2023; Hamidani et al., 2024; Lenart et al., 2025) or that consider purely geometric effects (Oganesyan et al., 2020; Beniamini et al., 2015). However, to account for the full Type I GRB dataset, a magnetar engine is apparently required for at least for some GRBs.
@@
If a BNS merger remnant is not too massive to promptly collapse into a BH ($M_{\rm rem}\lesssim 1.3M_{\rm TOV}$), it may evolve through transient phases (Fig. 1, lower panel).
II.2.1 Hypermassive Neutron Stars